OpenStudy (mayaal):

Find a ⋅ b. a = 4i - 4j, b = 4i + 5j -4 <8, 1> <16, -20> -36

OpenStudy (mayaal):

@SolomonZelman @tkhunny

OpenStudy (tkhunny):

You are calculating the inner product. You should get a scalar. Let's see it!

OpenStudy (mayaal):

ok,thnx for replying:)

OpenStudy (mayaal):

what do i do now?

OpenStudy (tkhunny):

Question: Why do you have an inner product question without ever seeing it done?

OpenStudy (mayaal):

i saw it done but i still didn't get it.

OpenStudy (mayaal):

@chantz417

OpenStudy (mayaal):

cn u help me?

OpenStudy (anonymous):

im not quite sure on this one sorry

OpenStudy (mayaal):

that's fine:)

OpenStudy (tkhunny):

Okay, use the Inner Product Formula. \(<e,f>\cdot<c,d> = e\cdot c + f\cdot d\) You have a = 4i - 4j, b = 4i + 5j \(<4,-4>\cdot<4,5> = 4\cdot 4 + (-4)\cdot 5 = 16 - 20 = -4\) That's about it. You do some. Let's see what you get.

OpenStudy (mayaal):

thnx alot @tkhunny :)

OpenStudy (tkhunny):

No samples from you?

OpenStudy (mayaal):

what do u mean?

OpenStudy (tkhunny):

Show me this one: a = 4i - 3j, b = 2i - 2j What is \(a\cdot b\)

OpenStudy (mayaal):

<4,-3><2,-2> <8,6>

OpenStudy (tkhunny):

<8,6> is no good. It's still a vector. You should get the scalar 8+6 = 14. Good work. Pretty close.

OpenStudy (mayaal):

ok,and what does a scalar exactly mean?

OpenStudy (tkhunny):

It's a number with no units or any other encumbrance. 14 lbs -- Not a scalar. <3,4> - Not a scalar. {1, 3, 6} -- Not a scalar. 12 -- Scalar

OpenStudy (mayaal):

thnxagain.wuld u mind helping me on another question?

OpenStudy (tkhunny):

Please post on a new thread. Show your work!

OpenStudy (mayaal):

ok.